OpenStudy (anonymous):

Find the domain of \[f(x)=(8-3x)^{1/2}\]

OpenStudy (nenadmatematika):

solve 8-3x>=0 and you'll get the domain

OpenStudy (anonymous):

I did that but I still don't understand... I got \[(-\infty,\frac{8}{3}]\]

OpenStudy (nenadmatematika):

your answer is right :D

OpenStudy (anonymous):

Oh, lol... I just realized that when I put in a number greater than 8/3... domain is so confusing.. sigh

OpenStudy (anonymous):

Thanks so much for you help once again!

OpenStudy (anonymous):

What would I do if that was x^2?